# Finite Sample Analysis of Distributional TD Learning with Linear Function Approximation

### v1

@@ 3.2 Linear-Categorical Projected Bellman Equation
${\bm{G}}(r):=\begin{bmatrix}{\bm{g}}_{0}(r),&\cdots,&{\bm{g}}_{K-1}(r)\end{bmatrix}\in{\mathbb{R}}^{K\times K},\quad\tilde{{\bm{G}}}(r):={\bm{G}}(r)-\bm{1}_{K}^{\intercal}\otimes{\bm{g}}_{K}(r)\in{\mathbb{R}}^{K\times K}.$
In analogy to approximation results of $\|{\bm{V}}^{\pi}-{\bm{V}}_{{\bm{\psi}}^{\star}}\|_{\mu_{\pi}}$ and $\ell_{2,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}^{\pi,K})$, the following lemma answers how close $\bm{\eta}_{{\bm{w}}^{\star}}$ is to $\bm{\eta}^{\pi}$, whose proof can be found in Appendix C.5.
Proposition 3.2 (Approximation Error of $\bm{\eta}_{{\bm{w}}^{\star}}$).
Consider the $\mu_{\pi}$-weighted $1$-Wasserstein error $W_{1,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}}}\right):=({\mathbb{E}}_{s\sim\mu_{\pi}}[W_{1}^{2}({\eta}^{\pi}(s),{\eta}_{{\bm{w}}^{\star}}(s))])^{1/2}$, it holds that
$W_{1,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}}}\right)\leq$ $K^{-1}(1-\gamma)^{-3}+(1-\gamma)^{-2}\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{K}\bm{\eta}^{\pi},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}^{\pi}}\right),$
@@ 3.3 Update Scheme of Linear-Categorical TD Learning
${\bm{W}}_{t}{=}{\bm{W}}_{t{-}1}{-}\alpha{\bm{\phi}}(s_{t})\!\!\left[{{\bm{\phi}}(s_{t})^{\intercal}{\bm{W}}_{t{-}1}{-}{\bm{\phi}}(s_{t{+}1})^{\intercal}{\bm{W}}_{t{-}1}\tilde{{\bm{G}}}^{\intercal}(r_{t})-\frac{1}{K{+}1}\!\left({\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t}){-}\bm{1}_{K}}\right)^{\intercal}}\right],$ (9)
for any $t\geq 1$, where $\alpha$ is the constant step size. In this paper, we also consider the Polyak-Ruppert tail averaging $\bar{{\bm{w}}}_{T}:=\left({T/2+1}\right)^{-1}\sum_{t=T/2}^{T}{\bm{w}}_{t}$ as in the analysis of Linear-TD in (Samsonov et al., 2024b). Standard theory of LSA (Mou et al., 2020) tells us, under some conditions, if we take an appropriate constant step size, $\bar{{\bm{w}}}_{T}$ will converges to the solution ${\bm{w}}^{\star}$ with rate $T^{-1/2}$ as $T\to\infty$.
Remark 1 (Comparison with Existing Works).
Our Linear-CTD can be regarded as a preconditioned version of SSGD with the PMF representation considered in (Bellemare et al., 2019) and (Section 9.6 Bellemare et al., 2023). The preconditioning technique (Chen, 2005; Li, 2017) is a commonly used methodology to accelerate solving optimization problems by reducing condition number. We precondition the vanilla SSGD with the PMF representation by removing the matrix ${\bm{C}}^{\intercal}{\bm{C}}$ on the right of Eqn. (23), whose condition number scales with $K^{2}$ (Lemma E.2). By introducing the preconditioner $\left({{\bm{C}}^{\intercal}{\bm{C}}}\right)^{-1}$, our Linear-CTD (Eqn. (9)) can achieve a convergence rate independent of $K$, which the vanilla form cannot achieve (see Appendix D.6). It is worth noting that our Linear-CTD (Eqn. (9)) is equivalent to the stochastic semi-gradient descent (SSGD) with CDF representation, which was also considered in (Lyle et al., 2019). See Appendix C.6 for the CDF representation, and Appendix C.7 for a self-contained derivation of SSGD with different representations. We further comment that our Linear-CTD can guarantee that the total mass of return distributions always be $1$, while the algorithms proposed in (Lyle et al., 2019) cannot.
@@ 4 Non-Asymptotic Statistical Analysis
In the task of distributional policy evaluation, the quality of estimator $\bm{\eta}_{\bar{\bm{w}}_{T}}$ can be measured by $\mu_{\pi}$-weighted $1$-Wasserstein error $W_{1,\mu_{\pi}}(\bm{\eta}_{\bar{\bm{w}}_{T}},\bm{\eta}^{\pi})$. By triangle inequality, the error can be decomposed as approximation error and estimation error: $W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}_{\bar{\bm{w}}_{T}})\leq W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}})+W_{1,\mu_{\pi}}(\bm{\eta}_{{\bm{w}}^{\star}},\bm{\eta}_{\bar{\bm{w}}_{T}})$. Proposition 3.2 provides an upper bound for the approximation error $W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}_{{\bm{w}}^{\star}})$, so it suffices to control the estimation error ${\mathcal{L}}(\bar{\bm{w}}_{T}):=W_{1,\mu_{\pi}}(\bm{\eta}_{{\bm{w}}^{\star}},\bm{\eta}_{\bar{\bm{w}}_{T}})$.
In the following, we give non-asymptotic convergence rates of ${\mathcal{L}}(\bar{\bm{w}}_{T})$. We start from the generative model setting, i.e. in the $t$-th iteration, we collect samples $s_{t}\sim\mu_{\pi}(\cdot),a_{t}\sim\pi(\cdot|s_{t}),r_{t}\sim{\mathcal{P}}_{R}(\cdot|s_{t},a_{t}),s_{t}^{\prime}\sim P(\cdot|s_{t},a_{t})$ from the generative model (we need to replace $s_{t+1}$ with $s_{t}^{\prime}$ in Linear-CTD Eqn. (9)). We give $L^{p}$ and high-probability convergence results in this setting. Then, we move back to the Markovian setting (i.e.  using the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$ introduced in Section 2.1), and give high-probability convergence results.
To facilitate the comparison of our results with those of Linear-TD, in the following, we denote by ${\bm{\theta}}_{0}=\left({{\bm{C}}\otimes{\bm{I}}_{d}}\right){\bm{w}}_{0}$ (resp. ${\bm{\theta}}^{\star}=\left({{\bm{C}}\otimes{\bm{I}}_{d}}\right){\bm{w}}^{\star}$) the initial (resp. optimal) parameters under CDF representation, where ${\bm{C}}$ is defined in Eqn. (16). See Appendix C.6 for details of CDF representation.
@@ 4.1 $L^{2}$ Convergence
We first provide non-asymptotic convergence rates of ${\mathbb{E}}^{1/2}[({\mathcal{L}}(\bar{\bm{w}}_{T}))^{2}]$.
@@
Suppose $K\geq(1-\gamma)^{-1}$, $T\geq 2$, $\alpha\in(0,(1-\sqrt{\gamma})/76)$ and ${\bm{w}}_{0}\in{\mathbb{R}}^{dK}$ is the initialization, then it holds that
@@
${\mathbb{E}}^{1/2}[({\mathcal{L}}(\bar{\bm{w}}_{T}))^{2}]\lesssim$ $\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\sqrt{T}(1-\gamma)\sqrt{\lambda_{\min}}}\left({1+\sqrt{\frac{\alpha}{(1-\gamma)\lambda_{\min}}}}\right)+\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{T\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}$
@@
$+\frac{(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}}{T\sqrt{\alpha}(1-\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}{+}\frac{1}{\sqrt{(1{-}\gamma)\lambda_{\min}}}}\right)\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|.$
@@
To prove Theorem 4.1, we conduct a fine-grained analysis of the linear-categorical Bellman equation and apply the exponential stability argument (Theorem 1 Samsonov et al., 2024b), which is outlined in Section 5. And the $L^{p}$ ($p>2$) convergence results can be found in Theorem D.1. Theorem 4.1 implies that learning the distribution of the return is as easy as learning its expectation (value function) with linear function approximation. Rowland et al. (2024b); Peng et al. (2024) discovered this phenomenon in the tabular setting, and we extend it to the function approximation setting.
Remark 2.
@@
The only difference between our Theorem 4.1 and the $L^{2}$ convergence rate of classic Linear-TD (Theorem 3 Samsonov et al., 2024b) lies in replacing $\left\|{\bm{\psi}^{\star}}\right\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}$ (resp. $\left\|{\bm{\psi}_{0}{-}\bm{\psi}^{\star}}\right\|$) with $K^{-1/2}(1{-}\gamma)^{-1}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$ (resp. $K^{-1/2}(1{-}\gamma)^{-1}\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|$). We claim that the two pairs should be of the same order respectively. Note that $\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}},\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|$ are of the order $O(\sqrt{K})$ (also dependent on ${\bm{\phi}}$, we omit it for brevity) if ${\eta}_{{\bm{w}}^{\star}}(s),{\eta}_{{\bm{w}}_{0}}(s)$ are valid probability distributions for all $s\in{\mathcal{S}}$. This is because in this case, the vector of CDF ${\bm{F}}_{{\bm{\theta}}}(s)=({\bm{\theta}}(k)^{\intercal}{\bm{\phi}}(s){+}({k{+}1})/({K{+}1}))_{k=0}^{K{-}1}\in\left[{0,1}\right]^{K}$ for ${\bm{\theta}}\in\left\{{{\bm{\theta}}^{\star},{\bm{\theta}}_{0}}\right\}$. While, in classic Linear-TD (Section 2.2), a proper estimate $\bm{\psi}$ should satisfy $V_{\bm{\psi}}(s)={\bm{\psi}}^{\intercal}{\bm{\phi}}(s)\in\left[{0,(1{-}\gamma)^{-1}}\right]$. It is thus reasonable to consider the two pairs as being of the same order, respectively. Similar arguments also holds in other convergence results presented in this paper. Therefore, in this sense, our results match those of classic Linear-TD.
@@
Under the same conditions as Theorem 4.1, for any $\varepsilon>0$, suppose
$T\gtrsim$ $\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}}\left({1+\frac{\alpha}{(1-\gamma)\lambda_{\min}}}\right)+\frac{\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\varepsilon\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}$
$+\frac{1}{\alpha(1-\gamma)\lambda_{\min}}\left({\log\frac{\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|}{\varepsilon\sqrt{K}(1{-}\gamma)}+\log\left({\frac{1}{T\sqrt{\alpha}(1{-}\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}{+}\frac{1}{\sqrt{(1{-}\gamma)\lambda_{\min}}}}\right)}\right)}\right),$
then it holds that ${\mathbb{E}}^{1/2}[({\mathcal{L}}(\bar{\bm{w}}_{T}))^{2}]\leq\varepsilon$.
@@
If we take the largest possible instance-independent step size, i.e. $\alpha\simeq(1-\gamma)$, and consider $\varepsilon\in(0,1)$, we obtain the sample complexity bound
$T=\widetilde{O}\left({\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}^{2}}}\right).$ (10)
@@
$T=\widetilde{O}\left({\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}}}\right),$ (11)
when we consider small enough $\varepsilon=\tilde{O}\left({\sqrt{\lambda_{\min}}}\right)$, or equivalently, large enough total update steps
$T=\widetilde{\Omega}\left({\frac{\frac{1}{K(1-\gamma)^{2}}\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}+1}{(1-\gamma)^{2}\lambda_{\min}^{2}}}\right).$ (12)
These theoretical results match the recent results for classic Linear-TD with constant step size (Patil et al., 2023; Li et al., 2024b; Samsonov et al., 2024b). It is possible to break the sample size barrier (Eqn. (12)) as in classic Linear-TD by applying the variance-reduction techniques (Li et al., 2023a) , we leave it as a future work.

### v2

@@ 3 Linear-Categorical TD Learning
${\bm{G}}(r):=\begin{bmatrix}{\bm{g}}_{0}(r),\ldots,{\bm{g}}_{K-1}(r)\end{bmatrix}\in{\mathbb{R}}^{K{\times}K},\quad\tilde{{\bm{G}}}(r):={\bm{G}}(r)-\bm{1}_{K}^{\top}\otimes{\bm{g}}_{K}(r)\in{\mathbb{R}}^{K{\times}K}.$
In analogy to the approximation bounds of $\|{\bm{V}}^{\pi}-{\bm{V}}_{{\bm{\psi}}^{\star}}\|_{\mu_{\pi}}$ and $W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}^{\pi,K})$, the following lemma answers how close $\bm{\eta}_{{\bm{\theta}}^{\star}}$ is to $\bm{\eta}^{\pi}$, whose proof can be found in Appendix D.5.
Proposition 3.2 (Approximation Error of $\bm{\eta}_{{\bm{\theta}}^{\star}}$).
It holds that
$W_{1,\mu_{\pi}}^{2}\left({\bm{\eta}^{\pi},\bm{\eta}_{{\bm{\theta}}^{\star}}}\right)\leq K^{-1}(1-\gamma)^{-3}+(1-\gamma)^{-2}\ell_{2,\mu_{\pi}}^{2}\left({{\bm{\Pi}}_{K}\bm{\eta}^{\pi},{\bm{\Pi}}_{{\bm{\phi}},K}^{\pi}\bm{\eta}^{\pi}}\right),$
@@
$\text{{{{Linear-CTD}}}: }\qquad{\bm{\Theta}}_{t}{\leftarrow}$ ${\bm{\Theta}}_{t{-}1}{-}\alpha{\bm{\phi}}(s_{t})\Big{[}{\bm{\phi}}(s_{t})^{\top}{\bm{\Theta}}_{t{-}1}{-}{\bm{\phi}}(s_{t{+}1})^{\top}{\bm{\Theta}}_{t{-}1}({\bm{C}}\tilde{{\bm{G}}}(r_{t}){\bm{C}}^{-1})^{\top}$ (13)
${-}\left({K{+}1}\right)^{-1}(\textstyle\sum_{j=0}^{K}{\bm{g}}_{j}(r_{t}){-}\bm{1}_{K})^{\top}{\bm{C}}^{\top}\Big{]},$
for any $t\geq 1$, where $\alpha$ is the constant step size. In Appendix F, we empirically validate the convergence of Linear-CTD through numerical experiments. It is easy to verify that, in the special case of the tabular setting (${\bm{\phi}}(s)=(\mathds{1}\{s{=}\tilde{s}\})_{\tilde{s}\in{\mathcal{S}}}$), Linear-CTD is equivalent to tabular categorical TD learning (Eqn. (9)). In this paper, we consider the Polyak-Ruppert tail averaging $\bar{{\bm{\theta}}}_{T}:=(T{/}2{+}1)^{-1}\sum_{t{=}T{/}2}^{T}{\bm{\theta}}_{t}$ (we use an even number $T$) as in the analysis of Linear-TD in Samsonov et al. [54]. Standard theory of LSA [37] says under some conditions, if we take an appropriate step size $\alpha$, $\bar{{\bm{\theta}}}_{T}$ will converge to the solution ${\bm{\theta}}^{\star}$ with rate $T^{-1/2}$ as $T\to\infty$.
Remark 1 (Comparison with Existing Linear Distributional TD Learning Algorithms).
Our Linear-CTD can be regarded as a preconditioned version of vanilla stochastic semi-gradient descent (SSGD) with the probability mass function (PMF) representation [4, Section 9.6]. See Appendix D.6 for the PMF representation, and Appendix D.7 for a self-contained derivation of SSGD with PMF representations. The preconditioning technique is a commonly used methodology to accelerate solving optimization problems by reducing the condition number. We precondition the vanilla algorithm by removing the matrix ${\bm{C}}^{\top}{\bm{C}}$ (see Eqn. (26)), whose condition number scales with $K^{2}$ (Lemma G.2). By introducing the preconditioner $({\bm{C}}^{\top}{\bm{C}})^{-1}$, our Linear-CTD (Eqn. (13)) can achieve a convergence rate independent of $K$, which the vanilla form cannot achieve. See Remark 5 and Appendix F for theoretical and experimental evidence respectively.
We comment that Linear-CTD (Eqn. (13)) is equivalent to SSGD with CDF representation, which was also considered in Lyle et al. [32]. The difference is that our Linear-CTD normalizes the distribution so that the total mass of return distributions always be $1$, while the algorithm in Lyle et al. [32] does not. See Appendix D.7 for a self-contained derivation of SSGD with CDF representations.
Bellemare et al. [4, Section 9.5] also proposed a softmax-based linear-categorical algorithm, which is closer to the practical C51 algorithm [2]. However, the nonlinearity due to softmax makes it difficult for analysis. We will leave the analysis for it as future work.
Remark 2 (Linear-CTD is mean-preserving).
A key property of Linear-CTD is mean preservation. That is, if we use identical initializations (${\mathbb{E}}_{G\sim\bm{\eta}_{{\bm{\theta}}_{0}}}[G]={\bm{V}}_{{\bm{\psi}}_{0}}$) and an identical data stream to update in both Linear-CTD and Linear-TD, it follows that ${\mathbb{E}}_{G\sim\bm{\eta}_{{\bm{\theta}}_{t}}}[G]={\bm{V}}_{{\bm{\psi}}_{t}}$ for all $t$. However, the mean-preserving property does not hold for the SSGD with the PMF representation. In this sense, our Linear-CTD is indeed the generalization of Linear-TD. See Appendix D.8 for details.
@@ 4 Non-Asymptotic Statistical Analysis
In our task, the quality of estimator $\bm{\eta}_{\bar{\bm{\theta}}_{T}}$ is measured by $\mu_{\pi}$-weighted $1$-Wasserstein error $W_{1,\mu_{\pi}}(\bm{\eta}_{\bar{\bm{\theta}}_{T}},\bm{\eta}^{\pi})$. By triangle inequality, the error can be decomposed into the approximation error and the estimation error: $W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}_{\bar{\bm{\theta}}_{T}}){\leq}W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}_{{\bm{\theta}}^{\star}}){+}W_{1,\mu_{\pi}}(\bm{\eta}_{{\bm{\theta}}^{\star}},\bm{\eta}_{\bar{\bm{\theta}}_{T}})$. Proposition 3.2 already provided an upper bound for the approximation error $W_{1,\mu_{\pi}}(\bm{\eta}^{\pi},\bm{\eta}_{{\bm{\theta}}^{\star}})$, so it suffices to control the estimation error $W_{1,\mu_{\pi}}(\bm{\eta}_{{\bm{\theta}}^{\star}},\bm{\eta}_{\bar{\bm{\theta}}_{T}})$, denoted ${\mathcal{L}}(\bar{\bm{\theta}}_{T})$.
In the following theorem, we give non-asymptotic convergence rates of ${\mathcal{L}}(\bar{\bm{\theta}}_{T})$. We start from the generative model setting, i.e., in the $t$-th iteration, we collect samples $s_{t}\sim\mu_{\pi}(\cdot),a_{t}\sim\pi(\cdot|s_{t}),(r_{t},s_{t}^{\prime})\sim{\mathcal{P}}(\cdot,\cdot|s_{t},a_{t})$ from the generative model, and we replace $s_{t+1}$ with $s_{t}^{\prime}$ in Eqn. (13). We give $L^{p}$ and high-probability convergence results in this setting. These results can be extended to the Markovian setting, i.e., using the streaming data $\left\{{\left({s_{t},a_{t},r_{t}}\right)}\right\}_{t=0}^{\infty}$
@@ 4.1 $L^{2}$ Convergence
We first provide non-asymptotic convergence rates of ${\mathbb{E}}^{1/2}[({\mathcal{L}}(\bar{\bm{\theta}}_{T}))^{2}]$, which do not grow with the number of supports $K$. The $L^{p}$ ($p>2$) convergence results can be found in Theorem E.1.
@@
For any $K\geq(1-\gamma)^{-1}$ and $\alpha\in(0,(1-\sqrt{\gamma})/76)$, it holds that
@@
${\mathbb{E}}^{1/2}[({\mathcal{L}}(\bar{\bm{\theta}}_{T}))^{2}]\lesssim$ $\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{V_{1}}+1}{\sqrt{T}(1-\gamma)\sqrt{\lambda_{\min}}}\left({1+\sqrt{\frac{\alpha}{(1-\gamma)\lambda_{\min}}}}\right)+\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{V_{1}}+1}{T\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}$
@@
$+\frac{(1-\frac{1}{2}\alpha(1-\sqrt{\gamma})\lambda_{\min})^{T/2}}{T\sqrt{\alpha}(1-\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}{+}\frac{1}{\sqrt{(1{-}\gamma)\lambda_{\min}}}}\right)\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|_{V_{2}},$
where $\left\|{{\bm{\theta}}^{\star}}\right\|_{V_{1}}:=\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}^{\star}}\right\|_{{\bm{I}}_{K}\otimes{\bm{\Sigma}}_{{\bm{\phi}}}}$ and $\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|_{V_{2}}:=\frac{1}{\sqrt{K}(1-\gamma)}\left\|{{\bm{\theta}}_{0}-{\bm{\theta}}^{\star}}\right\|$.
@@
In the upper bound, the first term of order $T^{-1{/}2}$ is dominant. The second term of order $T^{-1}$ has a worse dependence on $\alpha^{-1}$, leading to a sample size barrier (Eqn. (15)). The third term, corresponding to the initialization error, decays at a geometric rate and can be thus ignored. To prove Theorem 4.1, we conduct a fine-grained analysis of the linear-categorical Bellman equation and apply the exponential stability argument [54]. We outline the proof in Section 5. In Remark 3, we compare our Theorem 4.1 with the $L^{2}$ convergence rate of classic Linear-TD and conclude that learning the distribution of the return is as easy as learning its expectation (value function) with linear function approximation. Rowland et al. [51], Peng et al. [43] discovered this phenomenon in the tabular setting, and we extended it to the function approximation setting.
Remark 3 (Comparison with Convergence Rate of Linear-TD).
@@
The only difference between our Theorem 4.1 and the tight $L^{2}$ convergence rate of classic Linear-TD (see Appendix C.2) lies in replacing $\|\bm{\psi}^{\star}\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}$ (resp. $\|\bm{\psi}_{0}{-}\bm{\psi}^{\star}\|$) in Linear-TD with $\|{\bm{\theta}}^{\star}\|_{V_{1}}$ (resp. $\|{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}\|_{V_{2}}$). We claim that the two pairs should be of the same order, respectively. Note that $\|{\bm{\theta}}^{\star}\|_{V_{1}}$ and $\|{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}\|_{V_{2}}$ are of order ${\mathcal{O}}((1{-}\gamma)^{-1})$ if ${\eta}_{{\bm{\theta}}^{\star}}(s)$ and ${\eta}_{{\bm{\theta}}_{0}}(s)$ are valid probability distributions for all $s\in{\mathcal{S}}$. This is because in this case, $F_{k}(s;{\bm{\theta}})={\bm{\phi}}(s)^{\top}{\bm{\theta}}(k){+}({k{+}1})/({K{+}1})\in[0,1]$ for ${\bm{\theta}}\in\left\{{{\bm{\theta}}^{\star},{\bm{\theta}}_{0}}\right\}$. While in Linear-TD, $\|{\bm{\psi}}^{\star}\|_{{\bm{\Sigma}}_{{\bm{\phi}}}}$ and $\|{\bm{\psi}}_{0}{-}{\bm{\psi}}^{\star}\|$ are also of order ${\mathcal{O}}((1{-}\gamma)^{-1})$ if $V_{{\bm{\psi}}}(s)={\bm{\phi}}(s)^{\top}{\bm{\psi}}\in[0,(1{-}\gamma)^{-1}]$ for all $s\in{\mathcal{S}}$ and ${\bm{\psi}}\in\{{\bm{\psi}}^{\star},{\bm{\psi}}_{0}\}$. It is thus reasonable to consider the two pairs with the same order, respectively. Similar arguments also hold in other convergence results presented in this paper. Therefore, in this sense, our results match those of Linear-TD.
@@
Under the same conditions as in Theorem 4.1, for any $\varepsilon>0$, suppose
$T\gtrsim$ $\frac{\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{V_{1}}+1}{\varepsilon^{2}(1-\gamma)^{2}\lambda_{\min}}\left({1+\frac{\alpha}{(1-\gamma)\lambda_{\min}}}\right)+\frac{\left\|{{\bm{\theta}}^{\star}}\right\|_{V_{1}}+1}{\varepsilon\sqrt{\alpha}(1-\gamma)^{\frac{3}{2}}\lambda_{\min}}$
$+\frac{1}{\alpha(1-\gamma)\lambda_{\min}}\left({\log\frac{\left\|{{\bm{\theta}}_{0}{-}{\bm{\theta}}^{\star}}\right\|_{V_{2}}}{\varepsilon}+\log\left({\frac{1}{T\sqrt{\alpha}(1{-}\gamma)\sqrt{\lambda_{\min}}}\left({\frac{1}{\sqrt{\alpha}}{+}\frac{1}{\sqrt{(1{-}\gamma)\lambda_{\min}}}}\right)}\right)}\right).$
Then it holds that ${\mathbb{E}}^{1/2}[({\mathcal{L}}(\bar{\bm{\theta}}_{T}))^{2}]\leq\varepsilon$.
@@
If we take the largest possible instance-independent step size, i.e., $\alpha\simeq(1-\gamma)$, and consider $\varepsilon\in(0,1)$, we obtain the sample complexity bound
$T=\widetilde{{\mathcal{O}}}\left({\varepsilon^{-2}(1-\gamma)^{-2}\lambda_{\min}^{-2}\left({\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{V_{1}}+1}\right)}\right).$ (14)
@@
$T=\widetilde{{\mathcal{O}}}\left({\left({\varepsilon^{-2}+{\lambda_{\min}^{-1}}}\right)(1-\gamma)^{-2}\lambda_{\min}^{-1}\left({\left\|{{\bm{\theta}}^{\star}}\right\|^{2}_{V_{1}}+1}\right)}\right).$ (15)
There is a sample size barrier in the bound, that is, the dependence on $\lambda_{\min}$ is the optimal $\lambda_{\min}^{-1}$ only when $\varepsilon{=}\tilde{{\mathcal{O}}}(\sqrt{\lambda_{\min}})$, or equivalently, we require a large enough (independent of $\varepsilon$) update steps $T$.
These results match the recent results for classic Linear-TD with a constant step size [26, 54]. It is possible to break the sample size barrier in Eqn. (15) as in Linear-TD by applying variance-reduction techniques [27]. We leave it for future work.